OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was 3 times his age. How old is Sam now? a. 20 b. 10 c. 30 d. 15

OpenStudy (anonymous):

20 years old i believe

OpenStudy (anonymous):

Let the age of Coach is : x And that of Sam is : y x = 2y And 10 years ago, Coach was : (x - 10) Sam was: (y - 10) So, x- 10 = 3(y - 10) x - 3y = -20 2y - 3y = -20 y = 20 So, Sam is now 20 years old..

OpenStudy (anonymous):

THANKYOU GUYS :D

OpenStudy (akashdeepdeb):

Okay let Sam's current age be = Sam Currently Coach's age = 2*Sam yrs. and Sam's age = Sam yrs. Now we go back to his past. They are telling us: When Sam started soccer 10 years ago, his coach was 3 times his age. So Sam, then was = Sam-10 yrs. And thus his coach, then was = 3(Sam-10) yrs. So now we need to get an equation: Coach currently is 2*Sam So 10 years ago Coach was 2*Sam -10 yrs. And we also have ONE MORE equation of coach, 10 years ago = 3(Sam-10) EQUATE THEM. 2*Sam - 10 = 3(Sam-10) 2*Sam - 10 = 3*Sam - 30 So Sam = 20 years old. :) Understood? :)

OpenStudy (anonymous):

20 is correct.